# Grounding or Guessing? Visual Signals for Detecting Hallucinations in Sign Language Translation

### v1

@@ 5 Conclusion
In this work, we propose grounding usage as a new dimension for hallucination detection. Grounding usage quantifies how much each generated token is supported by visual input, and when combined with text-based signals (confidence, perplexity, or entropy), it improves hallucination detection by distinguishing visually grounded from visually unsupported tokens. Our experiments show that incorporating grounding usage with established text-based baselines improves predictive performance. Our analysis further reveals a systematic imbalance: gloss-free models trained without explicit visual supervision tend to under-use the visual input, which explains their higher hallucination rates relative to gloss-based models. The latter benefit from intermediate gloss supervision, which enforces stronger visual grounding. In other words, hallucinations in gloss-free SLT largely arise when the model “guesses” based on linguistic priors instead of truly attending to visual evidence.
@@ F.2 CHAIR Computation and Language-Specific Preprocessing
For Chinese, we cannot rely on whitespace or inflectional morphology. Instead, we: (i) normalize the script with OpenCC (Traditional $\to$ Simplified) when needed; (ii) normalize spacing around Chinese punctuation to avoid spurious tokens; (iii) tokenize using jieba; and (iv) apply a Chinese stopword list to remove frequent grammatical particles and function words. Since Chinese lacks inflectional morphology, no stemming is applied. If spaCy’s pipeline is used, we additionally filter by PoS tags (same content set as above).

### v2

@@ 5 Conclusion
In this work, we propose a grounding usage based reliability measure as a new dimension for hallucination detection. Grounding usage quantifies how much each generated token is supported by visual input, and when combined with text-based signals (confidence, perplexity, or entropy), it improves hallucination detection by distinguishing visually grounded from visually unsupported tokens. Our experiments show that incorporating grounding usage with established text-based baselines improves predictive performance. Our analysis further reveals a systematic imbalance: gloss-free models trained without explicit visual supervision tend to under-use the visual input, which explains their higher hallucination rates relative to gloss-based models. The latter benefit from intermediate gloss supervision, which enforces stronger visual grounding. In other words, hallucinations in gloss-free SLT largely arise when the model “guesses” based on linguistic priors instead of truly attending to visual evidence.
@@ F.2 CHAIR Computation and Language-Specific Preprocessing
For Chinese, we cannot rely on whitespace or inflectional morphology. Instead, we: (i) normalize the script with OpenCC (Traditional $\to$ Simplified) when needed; (ii) normalize spacing around Chinese punctuation to avoid spurious tokens; (iii) tokenize using jieba; and (iv) apply a Chinese stopword list to remove frequent grammatical particles and function words. Since Chinese lacks inflectional morphology, no stemming is applied.